# Timelike entanglement entropy with gravitational anomalies

### v1

@@ Abstract
We study the timelike entanglement entropy (TEE) in two dimensional conformal field theories (CFT) with gravitational anomalies. We employ analytical continuation to compute the timelike entanglement entropy for a pure timelike interval in such CFTs. We find that, unlike the real part, the imaginary part of the TEE displays an asymmetric dependence on the central charges of the left and right moving modes. We propose that the asymmetric dependence on central charges of the imaginary part of the TEE can be used to probe the presence of gravitational anomalies in chiral CFT. Furthermore, we propose a holographic construction to obtain the timelike entanglement entropy from the bulk dual geometries involving topologically massive gravity in AdS3. The holographic results obtained matches exactly with the dual field theory results.
@@ 1 Introduction
The study of quantum entanglement in extended many body systems has recently become a crucial area of research yielding insights across condensed matter physics, quantum gravity and black hole physics. A key tool in this investigation is entanglement entropy which quantifies the entanglement in a given bipartite pure states. For two dimensional conformal field theories (CFTs), the replica technique provides a means to calculate this entanglement entropy [1, 2]. However when left and right moving central charges in a CFT become unequal indicating a gravitational anomaly [3], the CFT exhibits a breakdown of stress-energy conservation at the quantum level. This anomaly arising from the non conservation of the stress-energy tensor implies a coordinate dependent description of the theory. The authors [4] obtained the entanglement entropy in two dimensional conformal field theories with gravitational anomalies using the replica technique and showed that it has an additional contribution coming from the Lorentz anomaly. Alternative methods such as the generalized Rindler approach have also been employed to compute entanglement entropy in such CFTs [5]. Subsequent investigations have focused on analyzing other bipartite mixed state entanglement measures such as reflected entropy, entanglement negativity and balanced partial entanglement entropy (BPE) within these anomalous CFTs [6, 7].
@@
In the context of AdS/CFT, [4] provides a bulk description for the holographic entanglement entropy in the presence of a gravitational anomaly where the bulk dual is described by the topologically massive gravity (TMG) in an asymptotically AdS3 spacetime. The action for the TMG-AdS3 consists of usual Einstein-Hilbert term and a gravitational Chern-Simons term which alters the motion of massive spinning particles propagating in the bulk geometry. Their worldlines take the form of ribbons incorporating a normal frame at each point. Due to its explicit dependence on the connection, the Chern-Simons term is not manifestly invariant under diffeomorphism. This absence of invariance directly results in unequal central charges in the dual field theory. The contribution of the Chern-Simons term to entanglement entropy is determined by the boost needed to propagate the normal frame along the worldline. Consequently, the holographic entanglement entropy is given by the sum of the geodesic length (calculated using the RT formula [8, 9]) and the twist of the normal frame along the geodesic [4]. The holographic entanglement entropy can also be determined using the Rindler method as discussed in [10]. Recently, the correction to the holographic entanglement entropy (also BPE) has also been reproduced through the application of newly introduced surface termed inner RT (IRT) surface [11]. See [12, 13, 14, 15, 16, 17, 18, 19, 20, 21] for more development related to TMG-AdS3/CFT2 duality.
@@
Previous research on entanglement entropy has primarily focused on cases where the subsystem in question is a spacelike separated region. However, recent studies have extended this to include timelike scenarios. In this regard, the authors [22, 23] introduced a new entanglement measure for timelike subsystems termed timelike entanglement entropy. It can be defined as the analytic continuation of entanglement entropy of a spacelike subsystem to a timelike one and exhibits complex values in a CFT2. Holographically as shown in [22, 23], the timelike entanglement entropy (TEE) can be obtained by considering combined spacelike and timelike extremal surfaces (geodesics) in the bulk. Specifically, the spacelike extremal surface contributes the real part of TEE while the timelike extremal surface contributes the imaginary part which is obtained by embedding the Poincaré patch in the global patch. Moreover it has been argued that TEE is a specific example of pseudo entropy [24, 25, 26]. Since there can be many multiple combined spacelike and timelike surfaces homologous to the boundary subsystem which gives infinitely many different complex valued areas, [27] introduced the complex-valued weak extremal surface (CWES) criteria which is a complex valued generalization of the Ryu-Takayanagi extremal surface. The unique surface is then selected by choosing the CWES with the lowest imaginary area. On the other hand, for geometries like Lifshitz spacetime where a global embedding is not known, the holographic computation for TEE must be performed entirely within the Poincaré patch. As demonstrated in [28] by extremization of length functional in Lifshitz metric, the holographic TEE involves two spacelike geodesics connecting the timelike interval endpoints to infinity and a timelike geodesic linking the spacelike geodesic endpoints. The timelike geodesic is smoothly joined with the spacelike geodesics at infinity. This method also reproduces the TEE results in AdS3/CFT2 case [28]. Subsequently this approach was extended to holographic TEE in non-conformal theories and non-relativistic theories [29, 30]. In [31], the authors obtain the holographic TEE by suitably generalizing the Ryu-Takayanagi formula to a complexified bulk geometry (refer to [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61] for recent progress). Very recently, a novel generalization of the density matrix to capture correlations between timelike separated subsystems have been introduced [62] and entanglement measures for timelike separated subsystems known as entanglement in time was proposed. They further demonstrated that entanglement in time coincides with the TEE for relativistic quantum field theories. This equivalence also applies to the anomalous CFT which we study here.
@@
In this paper, we examine how gravitational anomalies affect TEE in two dimensional CFTs considering both field theory and holographic descriptions. In this context, we first compute TEE at zero temperature, finite temperature with angular potential and finite angular potential with zero temperature in a CFT2 with a gravitational anomaly. This involves the analytic continuation of the entanglement entropy of a spacelike interval to a timelike one. The resulting TEE is complex and reveals distinct central charge dependencies in its real and imaginary parts. Its imaginary part depends on a single central charge while the real part on both left and right central charges. We then compute TEE holographically, showing that it involves summing on-shell actions of a massive spinning particle moving on the spacelike and timelike extremal worldlines. Given that the holographic TEE involves both spacelike and timelike geodesics, and the Chern-Simons contribution for spacelike geodesics is known from [4], we here derive the corresponding contribution for timelike geodesics. The key difference is that the Chern-Simons contribution now quantifies the rotation of normal frames along the timelike geodesic whereas it measures the twist (or Lorentz boost) of normal vectors for spacelike geodesics. This contribution from the Chern-Simons term exhibits a complex nature similar to the length of a timelike geodesic. Since the intersection of spacelike and timelike geodesics lie deep in the bulk, we prescribe a specific boundary condition for normal vectors which is based on the time direction in the boundary. Further these boundary conditions are used to obtain the Chern-Simons (anomalous) contribution from each type of geodesic to the TEE. We show that interestingly the bulk holographic results agree exactly with the dual field theory results.
@@
The rest of the paper is organized as follows. In section 2 we briefly review about the entanglement entropy in a CFT2 with gravitational anomaly. In section 3 we compute the TEE of a pure timelike interval using analytical continuation in a CFT2 with a gravitational anomaly. Subsequently in section 4 after reviewing the holographic TMG-AdS3/CFT2 duality, we derive the anomaly contribution due to timelike geodesic and propose a holographic construction for the TEE. We then obtain the holographic TEE in a CFT2 dual to Poincaré AdS3, non-extremal rotating BTZ black hole and extremal rotating BTZ black hole. Finally in section 5 we present our summary and discussions.
@@ 2 Entanglement entropy in CFT2 with gravitational anomalies
We first briefly review about the two dimensional CFT with a gravitational anomaly [3, 63]. Such theories are described by two copies of the Virasoro algebra with two unequal central charges $c_{L}$ and $c_{R}$, and there is a breakdown of energy-momentum conservation at the quantum level. The anomaly appears as a gravitational anomaly where the stress tensor is symmetric but not conserved. This non-conservation is captured by the anomalous divergence of the stress tensor as [12]
@@
From a bulk perspective, gravitational anomalies arise because the gravitational Chern-Simons term is only invariant under diffeomorphisms up to a boundary term that leads to a non-zero divergence of the dual field theory stress tensor. Alternatively, the anomaly can also arise from a broken Lorentz symmetry which is anomalies under local frame rotations resulting in a conserved stress tensor but not symmetric. This makes the theory frame dependent. One can move between these two perspectives by incorporating a local counterterm into the CFT generating functional [63, 64].
@@ 2.1 Zero temperature
In this subsection we review the entanglement entropy computation for a single interval at zero and finite temperature in a CFT2 with a gravitational anomaly as discussed in [4]. Consider a boosted interval described by $A\equiv[z_{1},z_{2}]=[(x_{1},t_{1}),(x_{2},t_{2})]$ and $B=A^{c}$ represent the rest of the system. This configuration is depicted in fig. 1.
@@
The entanglement entropy for a generic spacelike interval $A$ can be expressed in terms of the two point twist field correlators as [1, 2]
@@
where $\epsilon$ is a UV cut-off. Now analytically continuing to the Lorentzian signature via $z=x-t$, $\bar{z}=x+t$ and using $z=Re^{i\theta}$, the entanglement entropy can be rewritten as follows [4]
@@
In the absence of an anomaly ($c_{L}=c_{R}$), the entanglement entropy (2.6) simplifies to the familiar entanglement entropy of a single interval as described in [1, 2].
@@ 2.2 Finite temperature and angular potential
Here the second term represents the contribution from gravitational anomaly. The above expression reduces to the entanglement entropy of an interval in a CFT2 at finite temperature with an angular potential [65] in the absence of an anomaly.
@@ 2.3 Finite angular potential and zero temperature
Consider a generic spacelike interval $A\equiv[w_{1},w_{2}]$ in a CFT2 at a finite angular potential and zero temperature. For this case one of the inverse temperatures i.e $\beta_{L}$ goes to infinity and the other one $\beta_{R}$ remains finite. The conformal transformation relating correlation functions on the complex plane to those on the twisted cylinder is given by [^1] Since CFT2’s with a conserved angular momentum at zero temperature are dual to the extremal rotating BTZ black holes, the corresponding conformal transformation can be obtained by taking the limit $r\to\infty$ in (4.53) [66].
@@
By utilizing the above conformal transformation, the entanglement entropy for an interval $A$ can be obtained by evaluating the two-point twist correlator on the twisted cylinder as [66]
@@
where $w_{12}=x_{12}+i\tau_{12}$ and $\bar{w}_{12}=x_{12}-i\tau_{12}$. As described in [66], the first part of the above equation is the entanglement entropy of an interval $A$ at zero temperature and the second part is the entanglement entropy at an effective temperature $r_{0}/\pi\ell$. This structure arises because the left moving modes of the CFT are in their ground state while the right moving modes are at an effective temperature known as the Frolov-Thorne temperature $T_{FT}=r_{0}/\pi\ell$ [67]. To see the anomaly contribution explicitly, upon rearranging the above expression and analytically continuing to the Lorentzian signature leads to the entanglement entropy of an interval at zero temperature with finite angular momentum as follows
@@ 3.1 Zero temperature
This result (3.5b) generalizes that of [22, 23] for the timelike entanglement entropy in non-anomalous CFT. In the presence of gravitational anomaly, the TEE acquires an additional contribution proportional to the anomaly coefficient $c_{L}-c_{R}$ (third term in (3.5a)) and a partial cancellation of the usual imaginary part of the TEE with the contribution from the gravitational anomaly occurs. This results in an asymmetric dependence of the TEE on the central charges, with the real part depends on both central charges and the imaginary part depends only on one of the two central charges. In particular, the TEE of the left moving mode is purely real. This later property is an useful feature that one may exploit to detect the presence of gravitational anomaly in a chiral CFT.

### v2

@@ Abstract
We study the timelike entanglement entropy (TEE) in two dimensional conformal field theories (CFT) with gravitational anomalies. We employ analytical continuation to compute the timelike entanglement entropy for a pure timelike interval in such CFTs. We find that, unlike the real part, the imaginary part of the TEE displays an asymmetric dependence on the central charges of the left and right moving modes. We propose that the asymmetric dependence on central charges of the imaginary part of the TEE can be used to probe the presence of gravitational anomalies in chiral CFT. Furthermore, we propose a holographic construction to obtain the timelike entanglement entropy from the bulk dual geometries involving topologically massive gravity in AdS3. The holographic results obtained match exactly with the dual field theory results.
@@ 1 Introduction
The study of quantum entanglement in extended many body systems has recently become a crucial area of research yielding insights across condensed matter physics, quantum gravity and black hole physics. A key tool in this investigation is entanglement entropy which quantifies the entanglement in a given bipartite pure states. For two dimensional conformal field theories (CFTs), the replica technique provides a means to calculate this entanglement entropy Calabrese:2004eu ; Calabrese:2009qy . However when left and right moving central charges in a CFT become unequal indicating a gravitational anomaly Alvarez-Gaume:1983ihn , the CFT exhibits a breakdown of stress-energy conservation at the quantum level. This anomaly arising from the non conservation of the stress-energy tensor implies a coordinate dependent description of the theory. The authors Castro:2014tta obtained the entanglement entropy in two dimensional conformal field theories with gravitational anomalies using the replica technique and showed that it has an additional contribution coming from the Lorentz anomaly. Alternative methods such as the generalized Rindler approach have also been employed to compute entanglement entropy in such CFTs He:2023cju . Subsequent investigations have focused on analyzing other bipartite mixed state entanglement measures such as reflected entropy, entanglement negativity and balanced partial entanglement entropy (BPE) within these anomalous CFTs Basu:2022nds ; Wen:2022jxr .
@@
In the context of AdS/CFT, Castro:2014tta provides a bulk description for the holographic entanglement entropy in the presence of a gravitational anomaly where the bulk dual is described by the topologically massive gravity (TMG) in an asymptotically AdS3 spacetime. The action for the TMG-AdS3 consists of usual Einstein-Hilbert term and a gravitational Chern-Simons term which alters the motion of massive spinning particles propagating in the bulk geometry. Their worldlines take the form of ribbons incorporating a normal frame at each point. Due to its explicit dependence on the connection, the Chern-Simons term is not manifestly invariant under diffeomorphism. This absence of invariance directly results in unequal central charges in the dual field theory. The contribution of the Chern-Simons term to entanglement entropy is determined by the boost needed to propagate the normal frame along the worldline. Consequently, the holographic entanglement entropy is given by the sum of the geodesic length (calculated using the RT formula Ryu:2006bv ; Ryu:2006ef ) and the twist of the normal frame along the geodesic Castro:2014tta . The holographic entanglement entropy can also be determined using the Rindler method as discussed in Jiang:2019qvd . Recently, the correction to the holographic entanglement entropy (also BPE) has also been reproduced through the application of newly introduced surface termed inner RT (IRT) surface Wen:2024muv . See Kraus:2005zm ; Song:2016gtd ; Jiang:2017ecm ; Hijano:2017eii ; Wen:2018mev ; Gao:2019vcc ; Song:2019txa ; Sun:2008uf ; Basu:2021axf ; Wang:2023jfr for more development related to TMG-AdS3/CFT2 duality.
@@
Previous research on entanglement entropy has primarily focused on cases where the subsystem in question is a spacelike separated region. However, recent studies have extended this to include timelike scenarios. In this regard, the authors Doi:2022iyj ; Doi:2023zaf introduced a new entanglement measure for timelike subsystems termed timelike entanglement entropy. It can be defined as the analytic continuation of entanglement entropy of a spacelike subsystem to a timelike one and exhibits complex values in a CFT2. Holographically as shown in Doi:2022iyj ; Doi:2023zaf , the timelike entanglement entropy (TEE) can be obtained by considering combined spacelike and timelike extremal surfaces (geodesics) in the bulk. Specifically, the spacelike extremal surface contributes the real part of TEE while the timelike extremal surface contributes the imaginary part which is obtained by embedding the Poincaré patch in the global patch. Moreover it has been argued that TEE is a specific example of pseudo entropy Nakata:2020luh ; Mollabashi:2020yie ; Mollabashi:2021xsd . The pseudo entropy is defined by replacing the reduced density matrix with a non-Hermitian transition matrix which generally results in a complex-valued entropy. Furthermore, the holographic pseudo entropy in dS/CFT has been shown to be related to the holographic TEE in AdS/CFT via a double Wick rotation as discussed in Doi:2022iyj . Since there can be many multiple combined spacelike and timelike surfaces homologous to the boundary subsystem which gives infinitely many different complex valued areas, Li:2022tsv introduced the complex-valued weak extremal surface (CWES) criteria which is a complex valued generalization of the Ryu-Takayanagi extremal surface. The unique surface is then selected by choosing the CWES with the lowest imaginary area. On the other hand, for geometries like Lifshitz spacetime where a global embedding is not known, the holographic computation for TEE must be performed entirely within the Poincaré patch. As demonstrated in Basak:2023otu by extremization of length functional in Lifshitz metric, the holographic TEE involves two spacelike geodesics connecting the timelike interval endpoints to infinity and a timelike geodesic linking the spacelike geodesic endpoints. The timelike geodesic is smoothly joined with the spacelike geodesics at infinity. This method also reproduces the TEE results in AdS3/CFT2 case Basak:2023otu . Subsequently this approach was extended to holographic TEE in non-conformal theories and non-relativistic theories Afrasiar:2024lsi ; Afrasiar:2024ldn . In Heller:2024whi , the authors obtain the holographic TEE by suitably generalizing the Ryu-Takayanagi formula to a complexified bulk geometry (refer to Wang:2018jva ; Liu:2022ugc ; Guo:2022jzs ; Narayan:2022afv ; He:2023eap ; Jiang:2023ffu ; Narayan:2023ebn ; Chu:2023zah ; Jiang:2023loq ; He:2023wko ; Chen:2023eic ; He:2023ubi ; Narayan:2023zen ; Shinmyo:2023eci ; Guo:2023tjv ; Diaz:2023npx ; Kanda:2023jyi ; He:2023syy ; Das:2023yyl ; Grieninger:2023knz ; Guo:2024lrr ; Basu:2024bal ; Guo:2024edr ; Anegawa:2024kdj ; Goswami:2024vfl ; Jena:2024tly ; Xu:2024yvf ; Roychowdhury:2025ukl ; Guo:2025pru ; Roychowdhury:2025aye for recent progress). Note that while the quantum mechanical interpretation of TEE is not yet understood properly but the AdS/CFT correspondence suggests a clearer perspective. Here the imaginary part of TEE is related to the area of a timelike extremal surface in the bulk. This implies that the imaginary part of TEE may be linked to the emergence of the time coordinate in holography, much like the entanglement entropy is associated with the emergence of spatial dimensions. Very recently, a novel generalization of the density matrix to capture correlations between timelike separated subsystems have been introduced Milekhin:2025ycm and entanglement measures for timelike separated subsystems known as entanglement in time was proposed. They further demonstrated that entanglement in time coincides with the TEE for relativistic quantum field theories. This equivalence also applies to the anomalous CFT which we study here.
@@
In this paper, we examine how gravitational anomalies affect TEE in two dimensional CFTs considering both field theory and holographic descriptions. In this context, we first compute TEE at zero temperature, finite temperature with angular potential and finite angular potential with zero temperature in a CFT2 with a gravitational anomaly. This involves the analytic continuation of the entanglement entropy of a spacelike interval to a timelike one. The resulting TEE is complex and reveals distinct central charge dependencies in its real and imaginary parts. Its imaginary part depends on a single central charge while the real part on both left and right central charges. We then compute TEE holographically, showing that it involves summing on-shell actions of a massive spinning particle moving on the spacelike and timelike extremal worldlines. Given that the holographic TEE involves both spacelike and timelike geodesics, and the Chern-Simons contribution for spacelike geodesics is known from Castro:2014tta , we here derive the corresponding contribution for timelike geodesics. The key difference is that the Chern-Simons contribution now quantifies the rotation of normal frames along the timelike geodesic whereas it measures the twist (or Lorentz boost) of normal vectors for spacelike geodesics. This contribution from the Chern-Simons term exhibits a complex nature similar to the length of a timelike geodesic. Since the intersection of spacelike and timelike geodesics lie deep in the bulk, we prescribe a specific boundary condition for normal vectors which is based on the time direction in the boundary. Further these boundary conditions are used to obtain the Chern-Simons (anomalous) contribution from each type of geodesic to the TEE. We show that interestingly the bulk holographic results agree exactly with the dual field theory results.
@@
The rest of the paper is organized as follows. In section˜2 we briefly review about the entanglement entropy in a CFT2 with gravitational anomaly. In section˜3 we compute the TEE of a pure timelike interval using analytical continuation in a CFT2 with a gravitational anomaly. Subsequently in section˜4 after reviewing the holographic TMG-AdS3/CFT2 duality, we derive the anomaly contribution due to timelike geodesic and propose a holographic construction for the TEE. We then obtain the holographic TEE in a CFT2 dual to Poincaré AdS3, non-extremal rotating BTZ black hole and extremal rotating BTZ black hole. Finally in section˜5 we present our summary and discussions.
@@ 2 Entanglement entropy in CFT2 with gravitational anomalies
We first briefly review about the two dimensional CFT with a gravitational anomaly Alvarez-Gaume:1983ihn ; Alvarez-Gaume:1984zlq . Such theories are described by two copies of the Virasoro algebra with two unequal central charges $c_{L}$ and $c_{R}$, and there is a breakdown of energy-momentum conservation at the quantum level. The anomaly appears as a gravitational anomaly where the stress tensor is symmetric but not conserved. This non-conservation is captured by the anomalous divergence of the stress tensor as Kraus:2005zm
@@
From a bulk perspective, gravitational anomalies arise because the gravitational Chern-Simons term is only invariant under diffeomorphisms up to a boundary term that leads to a non-zero divergence of the dual field theory stress tensor. Alternatively, the anomaly can also arise from a broken Lorentz symmetry which is anomalies under local frame rotations resulting in a conserved stress tensor but not symmetric. This makes the theory frame dependent. One can move between these two perspectives by incorporating a local counterterm into the CFT generating functional Alvarez-Gaume:1984zlq ; Bardeen:1984pm .
@@ 2.1 Zero temperature
In this subsection we review the entanglement entropy computation for a single interval at zero and finite temperature in a CFT2 with a gravitational anomaly as discussed in Castro:2014tta . Consider a boosted interval described by $A\equiv[z_{1},z_{2}]=[(x_{1},t_{1}),(x_{2},t_{2})]$ and $B=A^{c}$ represent the rest of the system. This configuration is depicted in fig.˜1.
@@
The entanglement entropy for a generic spacelike interval $A$ can be expressed in terms of the two point twist field correlators as Calabrese:2004eu ; Calabrese:2009qy
@@
where $\epsilon$ is a UV cut-off. Now analytically continuing to the Lorentzian signature via z=xtz=x-t, $\bar{z}=x+t$ and using $z=Re^{i\theta}$, the entanglement entropy can be rewritten as follows Castro:2014tta
@@
In the absence of an anomaly ($c_{L}=c_{R}$), the entanglement entropy (2.6) simplifies to the familiar entanglement entropy of a single interval as described in Calabrese:2004eu ; Calabrese:2009qy .
@@ 2.2 Finite temperature and angular potential
Here the second term represents the contribution from gravitational anomaly. The above expression reduces to the entanglement entropy of an interval in a CFT2 at finite temperature with an angular potential Hubeny:2007xt in the absence of an anomaly.
@@ 2.3 Finite angular potential and zero temperature
Consider a generic spacelike interval $A\equiv[w_{1},w_{2}]$ in a CFT2 at a finite angular potential and zero temperature. For this case one of the inverse temperatures i.e $\beta_{L}$ goes to infinity and the other one $\beta_{R}$ remains finite. The conformal transformation relating correlation functions on the complex plane to those on the twisted cylinder is given by [^1] Since CFT2’s with a conserved angular momentum at zero temperature are dual to the extremal rotating BTZ black holes, the corresponding conformal transformation can be obtained by taking the limit $r\to\infty$ in (4.54) Caputa:2013lfa .
@@
By utilizing the above conformal transformation, the entanglement entropy for an interval $A$ can be obtained by evaluating the two-point twist correlator on the twisted cylinder as Caputa:2013lfa
@@
where $w_{12}=x_{12}+i\tau_{12}$ and $\bar{w}_{12}=x_{12}-i\tau_{12}$. As described in Caputa:2013lfa , the first part of the above equation is the entanglement entropy of an interval $A$ at zero temperature and the second part is the entanglement entropy at an effective temperature $r_{0}/\pi\ell$. This structure arises because the left moving modes of the CFT are in their ground state while the right moving modes are at an effective temperature known as the Frolov-Thorne temperature $T_{FT}=r_{0}/\pi\ell$ PhysRevD.39.2125 . To see the anomaly contribution explicitly, upon rearranging the above expression and analytically continuing to the Lorentzian signature leads to the entanglement entropy of an interval at zero temperature with finite angular momentum as follows
@@ 3.1 Zero temperature
This result (3.5b) generalizes that of Doi:2022iyj ; Doi:2023zaf for the timelike entanglement entropy in non-anomalous CFT. In the presence of gravitational anomaly, the TEE acquires an additional contribution proportional to the anomaly coefficient $c_{L}-c_{R}$ (third term in (3.5a)) and a partial cancellation of the usual imaginary part of the TEE with the contribution from the gravitational anomaly occurs. This results in an asymmetric dependence of the TEE on the central charges, with the real part depends on both central charges and the imaginary part depends only on one of the two central charges. In particular, the TEE of the left moving mode is purely real. This later property is an useful feature that one may exploit to detect the presence of gravitational anomaly in a chiral CFT.